# Revisiting the maximum mass limit of strange stars in quadratic curvature-matter coupled gravity

### v2

@@ V Stability Analysis through radial variation of adiabatic index
To substantiate the proposed theory of gravity, we analyse the adiabatic index within the present parameter space. Adiabatic index $(\Gamma)$ represents the specific heat ratio that embodies the stiffness of an equation of state. For an isotropic stellar configuration, the adiabatic index is written in the following form:
$\Gamma=\Bigg[1+\frac{\rho}{p}\Bigg]\frac{dp}{d\rho},$ (12)
@@
where $\frac{dp}{d\rho}$ represents the square of the sound velocity. Notably, we have used the MIT bag model EoS here, hence, $\frac{dp}{d\rho}=\frac{1}{3}$. Now, Heintzmann and Hillebrandt [68] showed that, in the Newtonian framework of macroscopic stability analysis for NS models subjected to small radial perturbations, maintaining a real eigenfrequency for the fundamental radial oscillation mode in an isotropic matter configuration requires the adiabatic index to be greater than $\frac{4}{3}$. Now, considering the present framework, we have graphically illustrated the radial variation of adiabatic index in Figs. 5, 6, 7 and 8. It must be noted that since our approach here is model independent, we have obtained these variations from the solution of TOV equations coupled with MIT bag model EoS.
@@
Figure 5: Radial variation of adiabatic index for $B_{g}=57.55~\mathrm{MeV/fm^{3}}$, and different combinations of $\alpha$ and $\beta$. Fig. 5.: $\beta=-0.5$ with $\alpha=0,~2.5,~4.7$ (red, green, blue). Fig. 5.: $\beta=0$ with $\alpha=0,~2.5,~5$ (red, green, blue). Fig. 5.: $\beta=0$ with $\alpha=10,~30,~50$ (red, green, blue). Fig. 5.: $\beta=0.5$ with $\alpha=0,~2.5,~5$ (red, green, blue).
Figure 6: Radial variation of adiabatic index for $B_{g}=95.11~\mathrm{MeV/fm^{3}}$, and different combinations of $\alpha$ and $\beta$. Figure 6.: $\beta=-0.5$ with $\alpha=0,~2.5,~5$ (red, green, blue). Figure 6.: $\beta=0$ with $\alpha=0,~2.5,~5$ (red, green, blue). Figure 6.: $\beta=0$ with $\alpha=10,~30,~50$ (red, green, blue). Figure 6.: $\beta=0.5$ with $\alpha=0,~2.5,~5$ (red, green, blue).
Figure 7: Radial variation of adiabatic index for $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and different combinations of $\alpha$ and $\beta$. Figure 7: $\alpha=0$, with $\beta=-0.82,~0$, and $0.82$ (red, green, blue). Figure 7: $\alpha=1.5$ with $\beta=-0.61,~0$ and, $0.61$ (red, green, blue).
@@
From Figs. 5, 6, 7 and 8, we note that the condition $\Gamma>\frac{4}{3}\approx 1.33$ is well-satisfied in this framework, which further substantiates our proposed framework.
VI Conclusions
@@
There exists an interesting correlation between $\beta$ and $\alpha$. For a particular bag parameter, if $\alpha$ is increased, the range of $\beta$ becomes narrow. When the bag parameter is increased, the corresponding numerical values of $\beta$ surely increase, but the same characteristic narrowing of the $\beta$ range persists with increasing $\alpha$. For $B_{g}=57.55~\mathrm{MeV/fm^{3}}$, $\alpha>52$ results in $\beta\rightarrow 0$, whereas in case of $B_{g}=95.11~\mathrm{MeV/fm^{3}}$, for $\alpha>78$, $\beta\rightarrow 0$. Therefore, if such high curvature corrections are involved, the gravity-matter coupling becomes insignificant. However, under such criteria, the present formalism can achieve a maximum mass of $2.42~\mathrm{M_{\odot}}$ for $\beta=0$, $\alpha=52$ and $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and $3.11~\mathrm{M_{\odot}}$ for $\beta=0$, $\alpha=78.20$ and $B_{g}=95.11~\mathrm{MeV/fm^{3}}$. Hence, we may elucidate that with increasing bag parameter, the range of $\alpha$ increases in the framework of $f(R+\alpha R^{2})$ gravity, which significantly increases the maximum mass in this scenario. Based on this result, it may be inferred that the inclusion of higher order curvature and non-minimal matter coupling provides an extended framework, within which the lighter companion of GW190814 with a maximum mass of $59~\mathrm{M_{\odot}} [22$], may also be interpreted as a strange star.
Table 5: Comparative analysis of maximum mass, radius, and Buchdahl limits for compact stars in various modified gravity frameworks considering different EoS.
@@
From Table 5, it is observed that, compared to previously established gravitational theories, the present formalism offers greater flexibility in exploring the maximum mass–radius range of compact objects. This extended range arises primarily from the inclusion of higher-order curvature corrections and the non-minimal matter–geometry coupling. Furthermore, Figs. 5, 6, 7, and 8 confirm that the proposed framework satisfies the stability conditions, thereby strengthening the theoretical viability of the model. To reiterate, the entire formalism shows how the quadratic curvature with non-minimal matter coupling alters the maximum mass-radius limit of strange stars, and this feature places the work right in the intersection of modified gravity phenomenology and compact star astrophysics. The deviations from GR, that appear in the analysis may open several potential avenues for new research. For example, as a possible research line, one may extend this framework to construct the tidal parameters, which could yield valuable insights into the tidal deformabilities of compact stars. With the rapid progress in gravitational wave observations, it is may be possible to probe the tidal properties of binary mergers with unprecedented precision. Using this framework such parameters can be studied with a high degree of accuracy. Further, the shift in the mass-radius range may be useful to study the compact stars in a more precise manner using the NICER and radio timing constraints. Apart from these, the future research may delve into scrutinising the validity of energy conditions in this framework is also a great way to test the viability of this study. Finally, we may conclude that these findings demonstrate the potential of higher-order curvature-matter coupled gravity to accommodate heavier strange stars, offering a promising framework for probing the fundamental nature of dense matter and strong-field gravity in extreme astrophysical environments.
@@ Acknowledgments
DB is thankful to the Department of Science and Technology (DST), Govt. of India, for providing the fellowship vide no: DST/INSPIRE Fellowship/2021/IF210761. PKC gratefully acknowledges support from IUCAA, Pune, India under Visiting Associateship programme. DB is also grateful to IUCAA, Pune, India for providing the visitor facilities. DB and PKC gratefully acknowledge the facilities provided by IUCAA during their visit, where this work was completed. The work of KB was supported by the JSPS KAKENHI Grant Number 24KF0100 and Competitive Research Funds for Fukushima University Faculty (25RK011).

### v3

@@ V.1 Radial variation of adiabatic index
To validate the use of a barotropic EoS, such as the MIT bag model EoS, within this proposed theory of gravity, we analyse the adiabatic index within the present parameter space. Adiabatic index $(\Gamma)$ represents the specific heat ratio that embodies the stiffness of an equation of state. For an isotropic stellar configuration, the adiabatic index is written in the following form:
$\Gamma=\left(1+\frac{\rho}{p}\right)\frac{dp}{d\rho},$ (12)
@@
where $\frac{dp}{d\rho}$ represents the square of the sound velocity. Notably, we have used the MIT bag model EoS here, hence, $\frac{dp}{d\rho}=\frac{1}{3}$. Now, Heintzmann and Hillebrandt [69] showed that, in the framework of macroscopic stability analysis for NS models subjected to small radial perturbations, maintaining a real eigenfrequency for the fundamental radial oscillation mode in an isotropic matter configuration requires the adiabatic index to be greater than $\frac{4}{3}$. Now, considering the present framework, we have graphically illustrated the radial variation of adiabatic index in Fig. 5. It must be noted that since our approach here is model independent, we have obtained these variations from the solution of TOV equations coupled with MIT bag model EoS.
@@
Figure 5: Radial variation of adiabatic index for (a) $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and (b) $B_{g}=95.11~\mathrm{MeV/fm^{3}}$. Here, $\beta=0.5$ and red, green, blue lines represent $\alpha=0,~2.5,$ and $5$, respectively.
@@
From Fig. 5, we note that the condition $\Gamma>\frac{4}{3}\approx 1.33$ is well-satisfied in this framework, which further substantiates our proposed framework.
V.2 Harrison-Zel’dovich-Novikov criterion
In literature, the Harrison-Zel’dovich-Novikov criterion [70] is a standard turning point analysis pertaining to the stability of stellar configuration and characterised by the computation of $\Big(\frac{dM}{d\rho_{c}}\Big)$, where $M$ is the maximum mass and $\rho_{c}$ is the central density of the stellar structure. The solution of TOV equations has provided central densities that lead up to the maximum mass. Any further increases in the central density will increase the maximum mass thereby increasing the gravitational pull leading to collapse. This feature is accurately described by the graphical representations of Figs. 6 and 7. In Fig. 6, we have plotted the mass vs central density, where the maximum mass points are denoted by the solid dots. Next, we have plotted $\Big(\frac{dM}{d\rho_{c}}\Big)$ vs $\rho_{c}$ in Fig. 7.
Figure 6: Mass-central density for (a) $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and (b) $B_{g}=95.11~\mathrm{MeV/fm^{3}}$. Here, $\beta=0.5$ and red, green, blue lines represent $\alpha=0,~2.5,$ and $5$, respectively.
Figure 7: $\frac{dM}{d\rho_{c}}$ vs $\rho_{c}$ for (a) $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and (b) $B_{g}=95.11~\mathrm{MeV/fm^{3}}$. Here, $\beta=0.5$ and red, green, blue lines represent $\alpha=0,~2.5,$ and $5$, respectively.
From Fig. 7, we note that the maximum mass points and the corresponding central density describe the onset of instability. Beyond these points $\Big(\frac{dM}{d\rho_{c}}\Big)<0$ describing a collapsed unstable configuration. To solidify this aspect, we have tabulated the parameters at the turning point in Table 5.
Table 5: Parameters obtained at the turning point.
$B_{g}~(\mathrm{MeV/fm^{3}})$ $\beta$ $\alpha$ At the turning point
$\rho_{c}~(\times 10^{15}~\mathrm{gm/cc})$ $\frac{dM}{d\rho_{c}}$
[…]
VI Detectability of strange stars and constraining the parameter space
The present $f(\tilde{R},T)=R+\alpha R^{2}+2\beta T$ framework leads to several distinctive astrophysical signatures that enable both the detectability of strange stars and direct constraints on the model parameter space. A key outcome of our analysis is the significant enhancement of the maximum stable mass, reaching values as high as $\sim 3.11~M_{\odot}$. Such large masses fall within the so-called mass gap between canonical neutron stars and low-mass black holes, thereby opening the possibility that some compact objects detected in GW events, most notably the secondary component of GW190814, may be strange stars rather than black holes. In particular, a compact object with mass $M\gtrsim 2.5~M_{\odot}$ and radius of order $\sim 10$ Km would be difficult to reconcile with standard hadronic equations of state in GR, while being naturally accommodated within the present model.
The increased compactness of strange star configurations in this theory also leaves a clear imprint on tidal interactions during binary inspirals. Specifically, the tidal Love numbers and the associated dimensionless tidal deformability $\Lambda(M)$ are systematically reduced relative to GR predictions. Consequently, joint constraints on masses and tidal deformabilities from LIGO-Virgo-KAGRA observations provide a powerful and quantitative test of the allowed $(\alpha,\beta)$ parameter space. Future high-precision measurements of inspiral waveforms will be particularly effective in discriminating between modified-gravity strange star models and standard NSs.
Beyond the inspiral phase, the raised maximum mass significantly affects post-merger dynamics. The delayed collapse of merger remnants implies the formation of longer-lived hypermassive or supramassive configurations, leading to characteristic shifts in the dominant post-merger gravitational-wave frequencies. Targeted searches for such spectral features in high signal-to-noise ratio detections may therefore provide an additional avenue to distinguish the present framework from GR-based expectations.
Electromagnetic observations further complement gravitational-wave probes. Strange stars, especially those with partially or fully bare quark surfaces, are expected to exhibit rapid neutrino cooling and distinctive thermal emission properties. These effects translate into characteristic X-ray cooling curves and spectral energy distributions that differ from those of hadronic NSs. Current and upcoming X-ray missions, such as NICER and XMM-Newton, can thus place meaningful constraints on the viability of the model.
Another promising observable lies in the moment of inertia and its relation to other macroscopic stellar properties. Owing to modifications in the internal structure induced by curvature-matter coupling, deviations from the universal I-Love-Q relations of GR are anticipated. Precise measurements of the moment of inertia in double pulsar systems would therefore offer a stringent test of the theory and help narrow down the admissible values of $\alpha$ and $\beta$.
Finally, pulsar timing observations provide complementary information through glitch activity. The presence or absence of a conventional crust in strange stars influences glitch amplitudes and recurrence statistics. Correlating glitch behaviour with independently inferred stellar masses and radii could offer indirect evidence in favour of-or against-the strange star interpretation within this modified gravity framework.
Taken together, a multi-messenger observational strategy combining mass-radius measurements (NICER), tidal deformability constraints (LIGO/Virgo/KAGRA), moment of inertia determinations (radio timing), and thermal emission studies (NICER/XMM-Newton) enables a comprehensive mapping of the allowed $(\alpha,\beta)$ parameter space. Such an approach also ensures consistency with existing solar-system and post-Newtonian bounds, thereby providing a robust pathway to validate or rule out the present theory.
VII Conclusions
@@
In this model, an interesting correlation exists between $\alpha$ and $\beta$. For a particular bag parameter, if $\alpha$ is increased, the range of $\beta$ becomes narrow. When the bag parameter is increased, the corresponding numerical values of $\beta$ increase, but the same characteristic narrowing of the $\beta$ range persists with increasing $\alpha$. For $B_{g}=57.55~\mathrm{MeV/fm^{3}}$, $\alpha>52$ results in $\beta\rightarrow 0$, whereas in case of $B_{g}=95.11~\mathrm{MeV/fm^{3}}$, for $\alpha>78$, $\beta\rightarrow 0$. Therefore, if such high curvature corrections are involved, the gravity-matter coupling becomes insignificant. However, under such criteria, the present formalism can achieve a maximum mass of $2.42~\mathrm{M_{\odot}}$ for $\beta=0$, $\alpha=52$ and $B_{g}=57.55~\mathrm{MeV/fm^{3}}$ and $3.11~\mathrm{M_{\odot}}$ for $\beta=0$, $\alpha=78.20$ and $B_{g}=95.11~\mathrm{MeV/fm^{3}}$. Hence, we may elucidate that with increasing bag parameter, the range of $\alpha$ increases in the framework of $f(R+\alpha R^{2})$ gravity, which significantly increases the maximum mass in this scenario. Based on this result, it may be inferred that the inclusion of higher order curvature and non-minimal matter coupling provides an extended framework, within which the lighter companion of GW190814 with a maximum mass of $59~\mathrm{M_{\odot}} [22$], may also be interpreted as a strange star.
The recent observations of Low-Mass X-ray Binaries using Rossi X-Ray Timing Explorer (RXTE) and XMM-Newton have provided the mass-radius of the NSs present in those binary systems. In addition, improvements in astronomical instrumentation have allowed us to constrain the mass-radius ranges of the lighter companions involved in GW events. To assess the validity of our theoretical model, we computed the radii of these NS and of the secondary GW-event objects by solving the TOV equations within our current framework. The resulting predictions are summarised in the table below:
Table 6: Radius prediction of recently observed NSs and GW events.
Compact objects Measured mass Measured radius $B_{g}$ $\beta$ $\alpha$ Predicted radius
from observations $\mathrm{(M_{\odot})}$ from observations $\mathrm{(Km)}$ $(\mathrm{MeV/fm^{3}})$ $\mathrm{(Km)}$
HER X-1 [71] $0.85^{+0.15}_{-0.15}$ $8.1^{+0.41}_{-0.41}$ 85 0.5 5 8.1
EXO 1745-248 [72] 1.4 11 60 -0.45 4.4 11
4U 1820-30 [73] $1.58^{+0.06}_{-0.06}$ $9.1^{+0.4}_{-0.4}$ 95.11 -0.5 0.15 9.1
4U 1608-52 [74] $1.74^{+0.14}_{-0.14}$ $9.3^{+1.0}_{-1.0}$ 95.11 0.05 43 9.3
GW170817 [75] 1.16-1.60 $10.7^{+2.1}_{-1.5}$ 57.55 0.3 1.1 $10.7^{+0.39}_{-0.38}$
GW190814 [22] $2.59^{+0.08}_{-0.09}$ 95.11 0.02 72 10.41
From Table 6, we note that the for a particular set of parameters, predicted and observed radii match very well. Hence, we may assert that the proposed framework may be deemed astrophysically relevant to study compact stellar structures.
Notably, compared to previously established gravitational theories, the present formalism offers greater flexibility in exploring the maximum mass-radius range of compact objects. This extended range arises primarily from the inclusion of higher-order curvature corrections and the non-minimal matter-geometry coupling. We have presented a comparative analysis among different gravitational theories in terms of mass, radius and Buchdahl limits for compact stars in Table 7.
Table 7: Comparative analysis of maximum mass, radius, and Buchdahl limits for compact stars in various modified gravity frameworks.
@@
Furthermore, Figs. 5, 6, and 7 confirm that the proposed framework satisfies the stability conditions, thereby strengthening the theoretical viability of the model. To reiterate, the entire formalism shows how the quadratic curvature with non-minimal matter coupling alters the maximum mass-radius limit of strange stars, and this feature places the work right in the intersection of modified gravity phenomenology and compact star astrophysics. The deviations from GR, that appear in the analysis may open several potential avenues for new research. For example, as a possible research line, one may extend this framework to construct the tidal parameters, which could yield valuable insights into the tidal deformabilities of compact stars. With the rapid progress in gravitational wave observations, it is may be possible to probe the tidal properties of binary mergers with unprecedented precision. Using this framework such parameters can be studied with a high degree of accuracy. Further, the wide mass-radius range may be useful to study the compact stars in a more precise manner using the NICER and radio timing constraints. Apart from these, the future research may delve into scrutinising the validity of energy conditions in this framework is also a great way to test the viability of this study. In this regard, we have summarised the possibilities of SS detection and constraining the present parameter space in Table 8.
Table 8: Observational signatures of strange stars in $f(\tilde{R},T)=R+\alpha R^{2}+2\beta T$ gravity and corresponding constraints on the model.
Observable Predicted signature Constrains
High-mass compact objects Stable configurations up to $\sim 3.11\,M_{\odot}$; mass-gap objects with $R\sim 10$ km $(\alpha,\beta)$ upper bounds
Tidal deformability Reduced $\Lambda(M)$ due to higher compactness GW inspiral posteriors
Post-merger GW spectrum Delayed collapse; shifted dominant frequencies Maximum mass, stiffness
Thermal emission and cooling Rapid neutrino cooling; distinct X-ray spectra Surface composition, $\beta$
Moment of inertia Deviations from GR I–Love–Q relations Internal structure, $\alpha$
Pulsar glitches Modified glitch statistics due to crust properties Strange star viability
Multi-messenger consistency Joint GW, X-ray, and radio constraints Allowed $(\alpha,\beta)$ region
Finally, we may conclude that these findings demonstrate the potential of higher-order curvature-matter coupled gravity to accommodate heavier strange stars, offering a promising framework for probing the fundamental nature of dense matter and strong-field gravity in extreme astrophysical environments.
@@ Acknowledgments
DB is thankful to the Department of Science and Technology (DST), Govt. of India, for providing the fellowship vide no: DST/INSPIRE Fellowship/2021/IF210761. PKC gratefully acknowledges support from IUCAA, Pune, India under Visiting Associateship programme. DB is also grateful to IUCAA, Pune, India for providing the visitor facilities. DB and PKC gratefully acknowledge the facilities provided by IUCAA during their visit, where this work was completed. The work of KB was partially supported by the JSPS KAKENHI Grant Numbers 24KF0100, 25KF0176 and Competitive Research Funds for Fukushima University Faculty (25RK011).